# Low-Rank Bandits via Tight Two-to-Infinity Singular Subspace Recovery

### v1

@@ 4 Singular Subspace Recovery
In the first phase of our algorithms for low-rank bandits, the singular subspaces of $M$ are estimated. In this section, we study the performance of this estimation procedure. We assume that the learner observes $T$ samples of noisy entries of $M$ chosen in an i.i.d. manner according to some distribution $\omega$ over $[m]\times[n]$, and that from these observations, she builds estimates of the singular subspaces of $M$. For simplicity, we assume that $\omega$ is known to the learner. This is without loss of generality in our settings [^4] Indeed, we can easily estimate the context distribution $\rho$ using only $\widetilde{O}(m)$ samples, and select $j$ uniformly at random.. We also assume that $\omega_{\min}=\min_{(i,j)\in[m]\times[n]}\omega_{i,j}>0$.
@@ 4.2 Guarantees in the two-to-infinity norm
Now, we present the main guarantee on these errors, which is adapted from (Stojanovic et al., 2023).
@@
$\epsilon_{\textup{Sub-Rec}}\lesssim\sqrt{\frac{\mu^{6}\kappa^{4}r^{3}(m+n)}{T\omega_{\min}mn(m\wedge n)}\log^{3}\left(\frac{(m+n)T}{\delta}\right)}$
@@
$T\gtrsim\frac{\mu^{4}\kappa^{2}r^{2}(m+n)}{\omega_{\min}mn}\log^{3}\left(\frac{(m+n)T}{\delta}\right).$
The proof of Theorem 4.1 is sketched in Appendix B. The analysis is technically involved and relies on combining the so-called leave-one-out argument (see (Chen et al., 2021) and references therein) together with a Poisson approximation technique (Stojanovic et al., 2023).
From Theorem 4.1, we see that if the matrix $M$ and the sampling distribution $\omega$ are homogeneous, then after only $\widetilde{\Omega}(m+n)$ observations, we can recover the singular subspaces with an error rate in the $\|\cdot\|_{2\to\infty}$ norm that scales as $\widetilde{O}(1/\sqrt{T})$ with high probability. The recovery error in $\|\cdot\|_{\textup{F}}$ would typically scale as $\widetilde{O}(\sqrt{(m+n)/T})$ (see Lemma B.7; it can also be established using classical arguments [^5] By classical arguments, we mean using a matrix concentration bound such as matrix Bernstein inequality and Davis-Kahan-Wedin Theorem to obtain a high probability bound on $\|\widehat{V}\widehat{V}^{\top}-VV^{\top}\|_{F}$. as in (Jun et al., 2019)). This suggests that the subspace recovery error (seen as a matrix) is delocalized, i.e., spread out along $m+n$ directions. Thus, the provided two-to-infinity norm guarantees offer a finer and more precise control over the subspace recovery error than the Frobenius norm guarantees. We take advantage of this when devising algorithms for contextual low-rank bandits.
@@ 5 Reduction to Misspecified Linear Bandits
In this section, we assume that we use $T_{1}$ reward observations to estimate the singular subspaces $U$ and $V$ of $M$ as described in the previous section. These $T_{1}$ first observations will constitute the first phase of our low-rank bandit algorithms. Based on the estimates of $U$ and $V$, one could, as in (Jun et al., 2019; Kang et al., 2022), reduce the problem faced by the learner in the remaining $(T-T_{1})$ rounds to an almost low-dimensional linear bandit; see Appendix H for a detailed discussion. The second phase of low-rank bandit algorithms would then just consist in applying an algorithm for such linear bandits. We explain in Appendix H that for regret minimization, this approach would however not lead to the tightest regret guarantees possible. Instead, we establish that the second phase of low-rank bandit algorithms can be reduced to a misspecified linear bandit with misspecification controlled using our subspace recovery error guarantees in the two-to-infinity norm. This new approach yields improved regret upper bounds, but also allows us to devise efficient algorithms for PE and BPI. In what follows, we make the reduction to a misspecified linear bandit precise.
@@
From (3), we can write for any $(i,j)\in[m]\times[n]$,
@@
With this interpretation, we obtain a misspecified linear bandit of dimension $d=r(m+n)-r^{2}$. The benefit of this reduction is that we can characterize the misspecification error $\max_{(i,j)\in[m]\times[n]}|\epsilon_{i,j}|$ using our subspace recovery guarantees in the two-to-infinity norm. Indeed, we establish:
@@
Let $\delta\in(0,1)$, $T_{1}>0$ be the number of samples observed by the sampling distribution $\omega$. Then, the event: for all $(i,j)\in[m]\times[n]$,
$|\epsilon_{i,j}|\lesssim\frac{\mu^{6}\kappa^{4}r^{3}(m+n)}{T_{1}\omega_{\min}\sqrt{mn}(m\wedge n)}\log^{3}\left(\frac{(m+n)T_{1}}{\delta}\right),$
@@
$T_{1}\gtrsim\frac{\mu^{4}\kappa^{2}r^{2}(m+n)}{\omega_{\min}mn}\log^{3}\left(\frac{(m+n)T_{1}}{\delta}\right).$
Corollary 5.1 follows immediately from Theorem 4.1. To see that, observe that we have: for all $(i,j)\in[m]\times[n]$,
@@
$\overset{(c)}{\leq}\epsilon_{\textup{Sub-Rec}}^{2}\sqrt{nm}\|M\|_{\max}$ (12)
where equality (a) follows from the fact that $\widehat{U}_{\perp}\widehat{U}_{\perp}^{\top}M=(I-\widehat{U}\widehat{U}^{\top})M$ since $\widehat{U}\widehat{U}^{\top}+\widehat{U}_{\perp}\widehat{U}_{\perp}^{\top}=I$, and similarly for $M\widehat{V}_{\perp}\widehat{V}_{\perp}^{\top}$, inequality (b) follows from $|x^{\top}My|\leq\|x\|_{2}\|y\|_{2}\|M\|_{\textup{op}}$ for any vectors $x\in\mathbb{R}^{m},y\in\mathbb{R}^{n}$, and finally inequality (c) follows from $\|M\|_{{\textup{op}}}\leq\sqrt{mn}\|M\|_{\max}$ for any $m\times n$ matrix.
We remark that when $M$ and $\omega$ are homogeneous, then $\max_{i,j\in[m]\times[n]}|\epsilon_{i,j}|=\widetilde{O}((m+n)/T_{1})$ which is due to the quadratic dependence on $\epsilon_{\textup{Sub-Rec}}$ appearing in (12).
@@ 6 Policy Evaluation
In this section, we present algorithms for PE and analyze their performance. The first algorithm, referred to as SIPS (Spectral Importance Propensity Score), consists of a single phase and exploits all samples to construct a rank-$r$ estimate of $M$. The value of the target policy is directly obtained using this estimate. We show that, when $M$ and $\omega$ are homogeneous, SIPS exhibits a minimax optimal sample complexity w.r.t. the matrix size $(m,n)$ and the accuracy level $\varepsilon$, but not w.r.t. the confidence level $\delta$. The design of the second algorithm, RS-PE (Recover Subspace for Policy Evaluation), leverages our two-phase approach. In its second phase, RS-PE refines the estimate of $M$ via a regularized least squares estimator applied to the misspecified linear bandit model obtained in the first phase. Applying this two-phase approach has significant benefits: (i) RS-PE enjoys instance-dependent sample complexity guarantees; (ii) When $M$ and $\omega$ are homogeneous, it is nearly minimax optimal w.r.t. the parameters $m,n$, $\varepsilon$ and $\delta$; (iii) as shown in Appendix J.1, it outperforms SIPS in numerical experiments. We conclude this section by deriving instance-dependent and minimax sample complexity lower bounds.
Throughout this section, we assume that the samples are gathered in an i.i.d. manner using a behavior policy $\pi^{b}$, and we wish to estimate the value $v^{\pi}$ of a target policy $\pi$. To simplify the notation, we abbreviate the (context, arm) distribution $\omega^{\pi^{b}}$ by $\omega$.
@@ 6.1 Spectral Importance Propensity Score
A natural approach to PE consists in directly constructing $\widehat{M}$, a rank-$r$ estimate of $M$ obtained via the spectral method described in (2), using all the available $T$ samples gathered under the (context, arm) distribution $\omega$. We estimate the value of the target policy $\pi$ by:
@@
Let $\varepsilon,\delta\in\left(0,1\right)$. For any target policy, $\hat{v}^{\pi}_{\text{\tt SIPS}}$ is $(\varepsilon,\delta)$-PAC as soon as
$T\gtrsim\frac{\mu^{11}\kappa^{4}r^{3}(m+n)^{3}}{\omega_{\min}mn(m\wedge n)^{2}\varepsilon^{2}}\log^{3}\!\left(\frac{(m+n)T}{\delta}\right)\!.$
The proof of this theorem, presented in Appendix D.1, relies on deriving entry-wise guarantees for the estimate $\widehat{M}$. These guarantees follow from the two-to-infinity guarantees provided in Theorem 4.1. Observe that when both $M$ and $\omega$ are homogeneous, the sample complexity derived in Theorem 6.1 scales as ${(m+n)\over\varepsilon^{2}}\log^{3}(1/\delta)$.
We note that a naive control by the Frobenius norm would not be sufficient to obtain an error guarantee that scales with $\sqrt{m+n}$. Indeed, by the Cauchy-Schwartz inequality, $|v^{\pi}-\hat{v}^{\pi}_{\text{\tt SIPS}}|\leq\|\omega^{\pi}\|_{F}\|M-\widehat{M}\|_{F}.$ If $\omega$ is homogeneous, $\|M-\widehat{M}\|_{F}$ typically scales with $\sqrt{mn\left(m+n\right)/T}$, and $\|\omega^{\pi}\|_{F}=\Theta\left(1/\sqrt{m}\right)$ when $\pi$ is deterministic. We consequently obtain $|v^{\pi}-\hat{v}^{\pi}_{\text{\tt SIPS}}|\lesssim\sqrt{n\left(m+n\right)/T}$, which does not improve over existing PE error bounds for contextual bandits with no structure (Yin & Wang, 2020).
@@ 6.2 Algorithm via the Two-phase Approach
$\widehat{\theta}=\left(\sum_{t=t_{1}}^{T}\phi_{i_{t},j_{t}}\phi_{i_{t},j_{t}}^{\top}+\tau I_{d}\right)^{-1}\left(\sum_{t=t_{1}}^{T}r_{t}\phi_{i_{t},j_{t}}\right),$ (14)
@@
$\hat{v}_{\text{\tt RS-PE}}^{\pi}:=\sum_{(i,j)\in[m]\times[n]}w^{\pi}_{i,j}\phi_{i,j}^{\top}\hat{\theta}.$ (15)
To analyze RS-PE, we introduce the following instance-dependent quantities. Define for all $(i,j)\in[m]\times[n]$,
@@
Let $\varepsilon,\delta\in(0,1)$. With the choices $T_{1}=\lfloor T/2\rfloor$ and $\tau\leq r(m\wedge n)^{-1}(\omega_{\min}/\omega_{\max})\log(16d/\delta)$, $\hat{v}^{\pi}_{\text{\tt RS-PE}}$ is $(\varepsilon,\delta)$-PAC as soon as
$T\gtrsim\frac{\|\psi_{\pi}\|^{2}_{2}}{\omega_{\min}\,\varepsilon^{2}}\log\left(\frac{e}{\delta}\right)+O\left({1\over\varepsilon^{4/3}}\right)\log^{3}\left(\frac{(m+n)T}{\delta}\right)\!,$
where the term $O({1\over\varepsilon^{4/3}})$ hides a multiplicative factor depending polynomially on the model parameters [^6] Refer to (36) in Appendix D.2 for a complete expression. $\mu,\kappa,r,m,n,(\omega_{\min}mn)^{-1}$.
@@
Asymptotically, when $\varepsilon$ tends to 0, these guarantees exhibit a better dependence in $\mu$, $\kappa$, $r$ and $\delta$ (and even in $(m,n)$ for non-homogeneous matrices $M$) compared to those of SIPS: for $\varepsilon$ small enough [^7] Refer to (40) in Appendix D.2 for a precise condition., the sample complexity of RS-PE scales as $\frac{\mu^{2}r(m+n)}{\omega_{\min}mn\varepsilon^{2}}\log(1/{\delta})$. When, in addition, $M$ and $\omega$ are homogeneous, this complexity scales as $\frac{(m+n)}{\varepsilon^{2}}\log(1/{\delta})$.
@@
We finally observe that without accounting for the low-rank structure, the minimax sample complexity of any PE algorithm would scale as ${mn\over\varepsilon^{2}}\log(1/\delta)$ (Yin & Wang, 2020). Exploiting the low-rank structure allows us to replace the factor $mn$ by $(m+n)$.

### v2

@@ 4 Singular Subspace Recovery
In the first phase of our algorithms for low-rank bandits, the singular subspaces of $M$ are estimated. In this section, we study the performance of this estimation procedure. We assume that the learner observes $T$ samples of noisy entries of $M$ chosen in an i.i.d. manner according to some distribution $\omega$ over $[m]\times[n]$, and that from these observations, she builds estimates of the singular subspaces of $M$. For simplicity, we assume that $\omega$ is known to the learner. This is without loss of generality in our settings [^5] Indeed, we can estimate the context distribution $\rho$ using only $\widetilde{O}(1/\rho_{\min})$ samples, and select $j$ uniformly at random (refer to Appendix I.1). We also assume that $\omega_{\min}=\min_{(i,j)\in[m]\times[n]}\omega_{i,j}>0$.
@@ 4.2 Guarantees in the Two-to-Infinity Norm
Now, we present the main guarantee on these errors, which is adapted from (Stojanovic et al., 2024).
@@
Let us define $\epsilon_{\textup{Sub-Rec}}:=\max(d_{2\to\infty}(U,\widehat{U}),d_{2\to\infty}(V,\widehat{V}))$ and $L:=\|M\|_{\max}\vee\sigma$. For any $\delta\in(0,1)$, the following event:
$\epsilon_{\textup{Sub-Rec}}\lesssim\sqrt{\frac{L^{2}\mu^{2}\kappa^{2}r(m+n)}{\sigma_{r}^{2}T\omega_{\min}(m\wedge n)}\log^{3}\left(\frac{(m+n)T}{\delta}\right)}$ (3)
@@
$T\gtrsim\frac{L^{2}(m+n)}{\sigma_{r}^{2}\omega_{\min}}\log^{3}\left(\frac{(m+n)T}{\delta}\right).$ (4)
The proof of Theorem 4.1 is sketched in Appendix B. The analysis is technically involved and relies on combining the so-called leave-one-out argument (see (Chen et al., 2021) and references therein) together with a Poisson approximation technique (Stojanovic et al., 2024).
From Theorem 4.1, we can immediately verify that if the matrix $M$ and the sampling distribution $\omega$ are homogeneous, then the subspace recovery error bound in the two-to-infinity norm is smaller by a factor of $\sqrt{m+n}$ than any achievable subspace recovery error bound in the Frobenius norm with our estimates of the singular subspaces. Indeed, assume additionally and only for simplicity that $\sigma\lesssim\|M\|_{\max}$, then after only $\widetilde{\Omega}(m+n)$ observations, we can recover the singular subspaces with an error rate in the $\|\cdot\|_{2\to\infty}$ norm that scales as $\widetilde{O}(1/\sqrt{T})$ with high probability. To see that, note that when $\sigma\lesssim\|M\|_{\max}$, then according to Lemma B.8, we have $L/\sigma_{r}=\Theta(1/\sqrt{mn})$. The recovery error in $\|\cdot\|_{\textup{F}}$ would typically scale as $\widetilde{O}(\sqrt{(m+n)/T})$ (see Lemma B.7 which follows from classical arguments as in (Jun et al., 2019)). This suggests that the subspace recovery error (seen as a matrix) is delocalized, i.e., spread out along $m+n$ directions. Thus, the provided two-to-infinity norm guarantees offer a finer and more precise control over the subspace recovery error than the Frobenius norm guarantees. We take advantage of this when devising algorithms for contextual low-rank bandits.
It is worth mentioning that the obtained error rate (3) in Theorem 4.1 exhibits a dependence on $L=\|M\|_{\max}\vee\sigma$ instead of only $\sigma$. This means that the proposed guarantees do not suggest exact subspace recovery when $\sigma\to 0$. This is to be expected because we rely on a spectral method that consists in truncating the SVD of $\widetilde{M}$. Indeed, such a method typically suffers from a bias that comes from setting unobserved entries to zero, as can be remarked in (1) (see also (Chen et al., 2021)). As we shall see in the coming sections, this affects the final bounds we obtain in all three learning tasks. Nonetheless, we provide a discussion in Appendix I.4 on how we may improve this dependence, which may be of independent interest. Indeed, one may resort to ideas based on nuclear norm penalization which have been shown recently to enjoy tight entry-wise matrix estimation guarantees (Chen et al., 2020). However, one has to take further care to ensure that the results based on such ideas would extend to our dependent noise setting. Finally, let us also mention that the dependence on $\sigma$, through $L$, in the sample complexity (4) provided in Theorem 4.1 is to be expected and cannot be further improved.
@@ 5 Reduction to Misspecified Linear Bandits
In this section, we assume that we use $T_{1}$ reward observations to estimate the singular subspaces $U$ and $V$ of $M$ as described in the previous section. These $T_{1}$ first observations will constitute the first phase of our low-rank bandit algorithms. Based on the estimates of $U$ and $V$, one could, as in (Jun et al., 2019; Kang et al., 2022), reduce the problem faced by the learner in the remaining $(T-T_{1})$ rounds to an almost low-dimensional linear bandit; see Appendix H for a detailed discussion. The second phase of low-rank bandit algorithms would then just consist in applying an algorithm for such linear bandits. We explain in Appendix H that for regret minimization, this approach would however not lead to the tightest regret guarantees possible. Instead, we establish that the second phase of low-rank bandit algorithms can be reduced to solving a misspecified linear bandit with misspecification controlled using our subspace recovery error guarantees in the two-to-infinity norm. This new approach yields improved regret upper bounds, but also allows us to devise efficient algorithms for PE and BPI. In what follows, we make the reduction to a misspecified linear bandit precise.
@@
From (5), we can write that for any $(i,j)\in[m]\times[n]$,
@@
With this interpretation, we obtain a misspecified linear bandit of dimension $d:=r(m+n)-r^{2}$. The benefit of this reduction is that we can characterize the misspecification $\epsilon_{\max}:=\max_{(i,j)\in[m]\times[n]}|\epsilon_{i,j}|$ using our subspace recovery guarantees in the two-to-infinity norm. Indeed, we establish:
@@
Let $\delta\in(0,1)$, $T_{1}>0$ be the number of samples observed by the sampling distribution $\omega$. Then, the event
$\epsilon_{\max}\lesssim\frac{L^{2}\mu^{2}\kappa^{3}r(m+n)}{\sigma_{r}T_{1}\omega_{\min}(m\wedge n)}\log^{3}\left(\frac{(m+n)T_{1}}{\delta}\right)$
@@
$T_{1}\gtrsim\frac{L^{2}(m+n)}{\sigma_{r}^{2}\omega_{\min}}\log^{3}\left(\frac{(m+n)T_{1}}{\delta}\right).$
Corollary 5.1 follows immediately from Theorem 4.1. To see that, observe that we have, for all $(i,j)\in[m]\times[n]$,
@@
$\overset{(c)}{\leq}\epsilon_{\textup{Sub-Rec}}^{2}\kappa\sigma_{r}(M)$ (8)
where equality (a) follows from the fact that $\widehat{U}_{\perp}\widehat{U}_{\perp}^{\top}M=(I-\widehat{U}\widehat{U}^{\top})M$ since $\widehat{U}\widehat{U}^{\top}+\widehat{U}_{\perp}\widehat{U}_{\perp}^{\top}=I$, and similarly for $M\widehat{V}_{\perp}\widehat{V}_{\perp}^{\top}$, inequality (b) follows from $|x^{\top}My|\leq\|x\|_{2}\|y\|_{2}\|M\|_{\textup{op}}$ for any vectors $x\in\mathbb{R}^{m},y\in\mathbb{R}^{n}$, and inequality (c) follows from the definition of the condition number $\kappa$.
We remark that when $M$ and $\omega$ are homogeneous, we have $\epsilon_{\max}=\widetilde{O}((m+n)/T_{1})$ due to the quadratic dependence on $\epsilon_{\textup{Sub-Rec}}$ appearing in (8).
@@ 6 Policy Evaluation
In this section, we present algorithms for PE and analyze their performance. The first algorithm, referred to as SIPS (Spectral Importance Propensity Score), consists of a single phase and exploits all samples to construct a rank-$r$ estimate of $M$. The value of the target policy is directly obtained using this estimate. We show that, when $M$ and $\omega$ are homogeneous, SIPS exhibits a minimax optimal sample complexity w.r.t. the matrix size $(m,n)$ and the accuracy level $\varepsilon$, but not w.r.t. the confidence level $\delta$. The design of the second algorithm, RS-PE (Recover Subspace for Policy Evaluation), leverages our two-phase approach. In its second phase, RS-PE refines the estimate of $M$ via a regularized least squares estimator applied to the misspecified linear bandit model obtained at the end of the first phase. Applying this two-phase approach has significant benefits: (i) RS-PE enjoys instance-dependent sample complexity guarantees; (ii) When $M$ and $\omega$ are homogeneous, it is nearly minimax optimal w.r.t. the parameters $m,n$, $\varepsilon$ and $\delta$; (iii) as shown in Appendix J.1, it outperforms SIPS in numerical experiments. We conclude this section by deriving instance-dependent and minimax sample complexity lower bounds.
Throughout this section, we assume that the observations are gathered in an i.i.d. manner using a behavior policy $\pi^{b}$, and we wish to estimate the value $v^{\pi}$ of a target policy $\pi$. To simplify the notation, we abbreviate the (context, arm) distribution $\omega^{\pi^{b}}$ by $\omega$.
@@ 6.1 Spectral Importance Propensity Score
A natural approach to PE consists in directly constructing $\widehat{M}$, a rank-$r$ estimate of $M$ obtained via the spectral method described in (2), using all of the $T$ available samples gathered under the (context, arm) distribution $\omega$. We estimate the value of the target policy $\pi$ by:
@@
Let $\varepsilon\in\left(0,\|M\|_{\max}\right)$ and $\delta\in\left(0,1\right)$. For any target policy $\pi$, $|v^{\pi}-\hat{v}^{\pi}_{\text{\tt SIPS}}|\leq\varepsilon$ with probability larger than $1-\delta$ as soon as
$T\gtrsim\frac{L^{2}\mu^{6}\kappa^{4}r^{3}(m+n)}{\omega_{\min}(m\wedge n)^{2}\varepsilon^{2}}\log^{3}\!\left(\frac{(m+n)T}{\delta}\right)\!.$
The proof of this theorem, presented in Appendix D.1, relies on deriving entry-wise guarantees for the estimate $\widehat{M}$. These guarantees follow from the two-to-infinity guarantees provided in Theorem 4.1. Observe that when both $M$ and $\omega$ are homogeneous, the sample complexity guarantees derived in Theorem 6.1 scale as ${L^{2}(m+n)\over\varepsilon^{2}}\log^{3}(1/\delta)$. The dependence in $L$ is a consequence of the bias induced by setting the unobserved entries of $M$ to 00 in the definition (1) of $\widetilde{M}$, as explained in Section 4.2.
Note that a naive control by the Frobenius norm would not be sufficient to obtain an error guarantee that scales with $\sqrt{m+n}$. Indeed, by the Cauchy-Schwartz inequality, $|v^{\pi}-\hat{v}^{\pi}_{\text{\tt SIPS}}|\leq\|\omega^{\pi}\|_{F}\|M-\widehat{M}\|_{F}.$ If $\omega$ is homogeneous, $\|M-\widehat{M}\|_{F}$ typically scales with $\sqrt{mn\left(m+n\right)/T}$, and $\|\omega^{\pi}\|_{F}=\Theta\left(1/\sqrt{m}\right)$ when $\pi$ is deterministic. We consequently obtain $|v^{\pi}-\hat{v}^{\pi}_{\text{\tt SIPS}}|\lesssim\sqrt{n\left(m+n\right)/T}$, which does not improve over existing PE error bounds for contextual bandits with no structure (Yin & Wang, 2020).
@@ 6.2 Algorithm via the Two-Phase Approach
$\hat{\theta}=\left(\sum_{t=t_{1}}^{T}\phi_{i_{t},j_{t}}\phi_{i_{t},j_{t}}^{\top}+\tau I_{d}\right)^{-1}\left(\sum_{t=t_{1}}^{T}r_{t}\phi_{i_{t},j_{t}}\right),$ (10)
@@
$\hat{v}_{\text{\tt RS-PE}}^{\pi}:=\sum_{(i,j)\in[m]\times[n]}\omega^{\pi}_{i,j}\phi_{i,j}^{\top}\hat{\theta}.$ (11)
To analyze RS-PE, we introduce the following instance-dependent quantities. Define, for all $(i,j)\in[m]\times[n]$,
@@
Let $\varepsilon\in\left(0,\|M\|_{\max}\right)$ and $\delta\in\left(0,1\right)$. With the choices $T_{1}=\lfloor T/2\rfloor$ and $\tau\leq r(m\wedge n)^{-1}(\omega_{\min}/\omega_{\max})\log(16d/\delta)$, for any target policy $\pi$, $|v^{\pi}-\hat{v}^{\pi}_{\text{\tt RS-PE}}|\leq\varepsilon$ with probability larger than $1-\delta$ as soon as
$T\gtrsim\frac{\sigma^{2}\|\psi_{\pi}\|^{2}_{2}}{\omega_{\min}\,\varepsilon^{2}}\log\left(\frac{e}{\delta}\right)+\frac{K}{\varepsilon^{4/3}}\log^{3}\!\left(\frac{(m+n)T}{\delta}\right)\!,$
where $K$ depends polynomially on the model parameters [^6] Refer to (34) in Appendix D.2 for a complete expression. $\|M\|_{\max},\sigma,\mu,\kappa,r,m,n,(\omega_{\min}mn)^{-1}$.
@@
Asymptotically, when $\varepsilon$ tends to 0, these guarantees exhibit a better dependence in $\mu$, $\kappa$, $r$, $\|M\|_{\max}$ and $\delta$ (and even in $(m,n)$ for non-homogeneous matrices $M$) compared to those of SIPS: for $\varepsilon$ small enough [^7] Refer to (38) in Appendix D.2 for a precise condition., the sample complexity guarantees of RS-PE scale as $\frac{\sigma^{2}\mu^{2}r(m+n)}{\omega_{\min}mn\varepsilon^{2}}\log(1/{\delta})$. When, in addition, $M$ and $\omega$ are homogeneous, they scale as $\frac{\sigma^{2}(m+n)}{\varepsilon^{2}}\log(1/{\delta})$.
@@
Observe further that in any case, the first term in the sample complexity guarantees of RS-PE scales with $\sigma^{2}$. In contrast, the guarantees of SIPS scale with $L^{2}=\left(\|M\|_{\max}\vee\sigma\right)^{2}$ (see Theorem 6.1). The remainder term, however, still depends on $L$ through the factor $K$.
We finally note that without accounting for the low-rank structure, the minimax sample complexity of any PE algorithm would scale as ${mn\over\varepsilon^{2}}\log(1/\delta)$ (Yin & Wang, 2020). Exploiting the low-rank structure allows us to replace the factor $mn$ by $(m+n)$.